OpenStudy (anonymous):

find the value of "n" 3^6 x 3^n divided by 3^4=3^5

OpenStudy (anonymous):

i know the answer is 3 but i need to know how to work it out as we aren't allowed calculators

jimthompson5910 (jim_thompson5910):

do you know the rules of exponents?

OpenStudy (anonymous):

yes i do @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so what are some of the rules you know

OpenStudy (anonymous):

i know all the rules, in working it out i got up to, 3^6n-4= 3^5 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

3^6 x 3^n wouldn't give you 3^(6n)

jimthompson5910 (jim_thompson5910):

it gives you 3^6 x 3^n = 3^(6+n)

OpenStudy (anonymous):

do you know how to work it out? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

3^6 x 3^n divided by 3^4=3^5 would turn into 3^(6+n) divided by 3^4=3^5 and then that becomes 3^(6+n-4) = 3^5

jimthompson5910 (jim_thompson5910):

see how I'm getting all that?

OpenStudy (anonymous):

yes, that is what i got up to, but om stuck now:/ @jim_thompson5910

jimthompson5910 (jim_thompson5910):

well the bases are equal to 3, so the exponents must be equal, therefore, 3^(6+n-4) = 3^5 leads to 6+n-4 = 5

jimthompson5910 (jim_thompson5910):

now solve for n

OpenStudy (anonymous):

i don't know how to solve for n @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to combine like terms in 6+n-4 = 5

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

do you see the like terms on the left side?

OpenStudy (anonymous):

noo..:/

jimthompson5910 (jim_thompson5910):

6 and -4 are the like terms because they don't have a variable on them so 6-4 = 2

jimthompson5910 (jim_thompson5910):

6+n-4 = 5 becomes n+2 = 5

OpenStudy (anonymous):

thats what i did but it says the answer is 3

jimthompson5910 (jim_thompson5910):

now move over that +2

OpenStudy (anonymous):

move it where

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):